OpenStudy (sasogeek):

A mule and a donkey were going to market, laden with wheat. The mule said, "if you give me one measure, I should carry twice as much as you, but if I give you 1 measure, we should have equal burdens." what were each of their burdens?

OpenStudy (anonymous):

God, I have no idea how to solve this. Maybe if I sat down for a while and worked on it...

OpenStudy (sasogeek):

ok well i'll go ahead and solve it then? anyone solving it behind scenes so that i'll wait a while or i should go ahead?

OpenStudy (anonymous):

i think i have it

OpenStudy (y2o2):

me too

OpenStudy (sasogeek):

lemme know what you got

OpenStudy (sasogeek):

brb

OpenStudy (anonymous):

in fact i know i have it because the answer checks. so i will keep quiet

OpenStudy (sasogeek):

keep quiet? why lol, just go for it :D at least i know Laura said she won't be able to solve it after looking at it, you might wanna help her out :)

OpenStudy (y2o2):

I think it's 7 and 5 !!

OpenStudy (sasogeek):

yup

OpenStudy (y2o2):

i made it like that: (x+1) = 2(y-1) and (x-1) = (y+1)

OpenStudy (sasogeek):

let donkey's burden = x let mule's burden = y when donkey gives mule 1 measure, donkey's new burden = x-1 , and mule's new burden = y+1, but mules new burden (y+1) is twice donkey's new burden 2(x-1) => y+1=2(x-1)................... equation 1 when mule gives donkey 1 measure, mule's new burden = y-1 and donkey's new burden = x+1 but when mule gives 1 measure, their new burdens are the same => y-1 = x+1..................... equation 2 solve the simultaneous equation :)

OpenStudy (anonymous):

yikes. i got 6 and 8

OpenStudy (sasogeek):

well issue resolved now and we're all happy :D i wish i could write that in latex but it'll take me forever, i'm so slow with latex :(

OpenStudy (anonymous):

I got it!

OpenStudy (sasogeek):

:D